Calculate and describe exactly how you would make the following solutions:

           

            20 ml of a 100 mg ml-1 solution of tetracycline in ddH2O.
Answer #1

The molecular formula of tetracyclin is C22H24N2O8

Its molar mass is = (22xAt.mass of C ) + (24x At,mass of H)+(2xAt.mass of N) +( 8x At.mass of O)

                          = (22x12 ) + ( 24x1) + ( 2x14 ) + (8 x16 )

                          = 444 g/mol

Given concentration is 100\frac{mg}{mL}=100\frac{g}{L}=100\frac{g}{L}\times \frac{1}{444}\frac{mol}{g}\times \frac{1}{L}=0.225 \frac{mol}{L}

Molarity , M=\frac{mass}{molarmass}\times \frac{1000}{volume in mL}

So mass of tetracyclin is mass = \frac{molarity \times molar mass \times volume in mL}{1000}

                                     mass = \frac{0.225 \times 444 \times 20}{1000}=1.998 g

So weigh exactly 1.998 g of etracyclin and transfer this into a 20 mL volumetric flask & add some amount of double distilled water & shake well.After complete dissolution of tetracyclin add water upto the mark( on the neck of volumetric flask).

Thus the required solution is prepared.